OpenStudy (anonymous):

Who can help me in this problems ? Please 1) W^2 + 18w + 77 2) x^2 - x - 42 3) 12d^2 + 4d - 1 4) 20x^2 + 22x - 12 5) 6g^2 + 11g -35 6) d^2 - 14d + 49 7) 49b^2 + 70b + 25 8) r^2 - 49 9) 4x^2 - 81y^2

OpenStudy (whpalmer4):

We can't help you if you don't tell us enough about what you are trying to do. Are you supposed to factor these, solve them with the quadratic formula, what? What have you tried doing so far?

OpenStudy (anonymous):

if you're factoring, then its (w+11)(w+7)=0 so your final answer is w=-11, -7

OpenStudy (anonymous):

for the first one at least... Is this for virtual school?

OpenStudy (anonymous):

Like factoring trinomials ? Ex: 6x^2 + 5x + 1 = 6x^2 + 3x + 2x+ 1 = (3x+1) (2x+1) Idk :/

OpenStudy (anonymous):

yeah, the first one factors... I haven't gotten to the others yet but they look like they all factor. #8 is actually a difference of squares so that ones really easy. what's the lesson supposed to be?

OpenStudy (anonymous):

Uff bro i really don't know ! so I want to help